OpenStudy (anonymous):

What’s the equation of a line that passes through points (0, -1) and (2, 3)?

OpenStudy (abb0t):

First, find the slope: \[slope = m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] then pick any of the two points into: \[(y-y_0)=m(x-x_0)\] Rearrange it in the form: y = mx+b

OpenStudy (anonymous):

Thanks! That was very helpful.